calculate the pH of [OH-] = 5.6x10-3M
Answer #1

Ans. Method 1:

            pOH = -log [OH-] = -log (5.6 x 10-3) = 2.25

Now,

                        pH + pOH = 14.00                           - equation 1

                        or, pH = 14.00 – pOH = 14.00 – 2.25 = 11.749 = 11.75

Thus, pH = 11.75

Method 2:                  [H+] [OH-] = 10-14                             - equation 2

            Or, [H+] = 10-14 / [OH-] = 10-14 / (5.6 x 10-3) = 1.7857 x 10-12

Now,

                        pH = -log [H+] = -log (1.7857 x 10-12) = 11.748 = 11.75

Thus, pH = 11.75